OpenStudy (piercetheveil47):

The distances in space are measured in light-years. The distance from Earth to a star is 5.9 × 10^13 kilometers. What is the distance, in light-years, from Earth to the star (1 light-year = 9.46 × 10^12 kilometers)? 1.61 light-years 3.56 light-years 5.58 light-years 6.23 light-years

OpenStudy (freckles):

you are converting one unit to another unit we have 1ly=9.45x10^12 km so we have you need to do: \[5.9 \times 10^{13} km \cdot \frac{1 ly}{9.46 \times 10^{12} km}\] that second fraction there is like multiplying by 1 since 1ly=9.45x10^12km

OpenStudy (freckles):

my anyways you will see the km measurements cancel

OpenStudy (freckles):

so you just need to divide the 5.9x10^13 by 9.46x10^12 to get the ly amount

OpenStudy (mathmath333):

^^ The distance from Earth to a star is \(\large\tt \begin{align} \color{black}{5.9 × 10^{13} ~~kilometers\\~\\ =5.9 × 10^{13} ~~kilometers\\~\\ = \dfrac{5.9 × 10^{13}}{9.46 × 10^{12}}~~light-years}\end{align}\)

OpenStudy (piercetheveil47):

should i divide 5.9 and 9.46??

OpenStudy (mathmath333):

yes also the exponent parts

OpenStudy (freckles):

\[\frac{5.9}{9.45} \cdot \frac{10^{13}}{10^{12}} ly\]

OpenStudy (piercetheveil47):

i got D??

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\checkmark}\end{align}\)

OpenStudy (piercetheveil47):

thank you both SO much!